OpenStudy (calculusxy):

Avery can run at 10 uph (units per hour). The bank of a river is represented by the line 4x + 3y = 12, and Avery is at (7, 5). How much time does Avery need to reach the river?

OpenStudy (calculusxy):

@satellite73

satellite73 (satellite73):

you are looking for the distance from \(4x+3y=12\) to \((7,5)\)

OpenStudy (calculusxy):

yes

OpenStudy (calculusxy):

should i set up a parametric?

satellite73 (satellite73):

dont think it is necessary

satellite73 (satellite73):

you could use calc if you want, that is one way

OpenStudy (calculusxy):

Do you mean use calculus? If so, I don't know much about it :(

satellite73 (satellite73):

oh then you cant use it you have to do something else

satellite73 (satellite73):

the slope of the line is what?

OpenStudy (calculusxy):

Can I find the line that is perpendicular to 4x+ 3y = 12 and also goes through the point (7, 5)? Then find the intersection of the two lines.

OpenStudy (calculusxy):

And after I find the intersection I can find the distance between (7, 5). But how do I find the time from there?

satellite73 (satellite73):

exactly

satellite73 (satellite73):

the time?

satellite73 (satellite73):

whatever it is time is distance divided by rate

OpenStudy (calculusxy):

It asks how much time it takes for her to reach the river

OpenStudy (calculusxy):

The rate would be the slope of the perpendicular line right?

satellite73 (satellite73):

that is a red herring, something to do at the end to keep busy

satellite73 (satellite73):

no the rate is teh rate

satellite73 (satellite73):

you are told the rate is 10

OpenStudy (calculusxy):

Oh I misread that. My bad.

satellite73 (satellite73):

find the distance, divide by 10 to find the time

OpenStudy (calculusxy):

OK I am doing that right now. Please give me a moment

satellite73 (satellite73):

ok

satellite73 (satellite73):

calculus may be easier...

OpenStudy (calculusxy):

Yeah but I'll get to that maybe in the next year or so :)

OpenStudy (calculusxy):

So far I am taking accelerated geometry

satellite73 (satellite73):

seems pretty speedy to me

OpenStudy (calculusxy):

Haha..

OpenStudy (mathmale):

Have you finished solving this problem? Hope so. If not, what have you left to do?

OpenStudy (calculusxy):

I get the Point of Intersection as (2.04, 1.28)

OpenStudy (calculusxy):

Is that correct?

satellite73 (satellite73):

ikd i didn't do it want me to check?

OpenStudy (calculusxy):

I would truly appreciate it if you can

satellite73 (satellite73):

ok give me a minute

OpenStudy (mathmale):

There is a formula for finding the distance from a point to a line; do you recall having seen such a formula before? I've forgotten which course you're in.

OpenStudy (calculusxy):

\(\large D = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\)

satellite73 (satellite73):

i get \[-3 x + 4 y =-1\] for the other line

OpenStudy (calculusxy):

Yes in slope-int form I get: y = (3/4)x - 1/4

OpenStudy (calculusxy):

I get my final answer to be 0.62 hours

satellite73 (satellite73):

that is what i got as well

satellite73 (satellite73):

distance \(6.2\) divide by 10

OpenStudy (calculusxy):

Thank you! Do you still have some more time to help me out?

satellite73 (satellite73):

sure